# Mitigating Propensity Bias of Large Language Models for Recommender Systems

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems are essential tools used on many platforms to help users navigate the vast amount of available information and find items that match their preferences (Gao et al., 2023; Wang et al., 2018; Wu et al., 2023). Due to the information overload problem (Zhang et al., 2017; He et al., 2020), collaborative filtering-based recommender systems have become widely used. The main research direction in this field focuses on improving the performance of these models by increasing the effectiveness of user and item representations (Zhang et al., 2019). To achieve this, many models have been proposed to mine useful information to enhance these representations. Graph neural networks play an important role in this endeavor (Wei et al., 2022; Wu et al., 2024). Recently, graph neural network-based recommender systems modeling user preferences based on historical interactions have achieved promising results (Wu et al., 2022a). Nevertheless, collaborative filtering models still face the problem of data sparsity (Cai et al., 2023; Yu et al., 2023a).
Researchers have argued that relying solely on interaction information may ignore other valuable data (Gao et al., 2023; Wang et al., 2020), such as rich information related to users and items in the real world. The lack of additional information can lead to a less informative representation learning. Previous research has demonstrated the importance of side information, also known as user and item descriptions, for collaborative filtering recommender systems to perform effective representation learning (Wei et al., 2022; Zhang et al., 2023b; Ren et al., 2024). However, these approaches tend to be based on domain-specific knowledge and are not sufficiently generalised for recommendation tasks that they do not know much about (Peng et al., 2023).
@@
Recently, Large Language Models (LLMs) equipped with powerful reasoning capabilities and extensive real-world knowledge have been widely applied across various domains, including recommender systems (Wei et al., 2024). These models, with their ability to process vast amounts of text data, offer new avenues for enhancing traditional recommender systems. However, directly deploying off-the-shelf LLMs for recommendation tasks presents several challenges (Wei et al., 2024; Ren et al., 2024; Zhang et al., 2024a; Lu et al., 2024). A key issue lies in the inherent mismatch between the generalisation capabilities of LLMs and the specific demands of personalised recommendation, particularly when it comes to capturing the nuanced preferences and behaviors of individual users. To bridge this gap, researchers have attempted to leverage LLMs to model human-like descriptions of users and items, thereby obtaining richer side information. The enriched data can then be aligned with collaborative representations learned from users’ historical interactions via techniques such as contrastive learning (Ren et al., 2024), representing a promising direction for integrating the reasoning capabilities of LLMs with the precision required for personalised recommendations.
However, the integration of LLMs may introduce an often overlooked challenge to recommender systems: the inherent biases in LLM training datasets (Jiang et al., 2024) inevitably introduce propensity bias. Specifically, LLM may be more inclined to generate responses that are biased toward certain cultures and viewpoints (Tao et al., 2023; Yu et al., 2023b; Sivaprasad et al., 2024), and the propensity bias would make the recommendations overgeneralise and ignore specific details. As shown in Figure 1, a user with ID 5596 in the Amazon dataset (He and McAuley, 2016) has previously read a range of thrilling and exciting books, yet the LLM overlooks this in its reasoning and generalises the user’s preference as simply “entertaining”.
@@
Furthermore, propensity bias not only impacts recommendation accuracy but also contributes to the homogenisation of LLM-generated representations (Barbero et al., 2024), which increases the risk of dimensional collapse. When using LLMs (Guo et al., 2024) or contrastive learning (Jing et al., 2022), feature representations are more likely to be restricted to a certain portion of the low-dimensional subspace, futher limiting their representational capability (Zhang et al., 2023c). As shown in Figure. 2, when using LightGCN as the backbone method to align LLMs for recommendation through contrastive learning, the model exhibits dimensional collapse when side information is encoded using MLP. This occurs because the side information repsentation generated by the LLM lacks structural information, making it difficult to align with collaborative information derived from historical interactions, thereby exacerbating the risk of dimensional collapse.
@@ 2.1. Bias in Artificial Intelligence
With the widespread use of artificial intelligence in real life, bias in artificial intelligence has attracted a great deal of attention (Zhang et al., 2023a; Ling et al., 2024). For example, Buolamwini and Gebru (Buolamwini and Gebru, 2018) showed that business gender classification systems used in various fields such as marketing, entertainment and healthcare have different representations of users’ skin colour and gender. Zhang et al. (Zhang et al., 2024c) demonstrated that the interaction structure in social networks can also contain preference information. It follows that artificial intelligence systems may display biases against specific groups embedded in the data or even amplify human biases. Natural Language Processing (NLP) systems are susceptible to bias as a result of the training corpus. Bolukbasi et al. (Bolukbasi et al., 2016) and Caliskan et al. (Caliskan et al., 2017) first found the presence of human-like preferences in word embedding models. Zhao et al. (Zhao et al., 2018) found a correlation between gender pronouns and stereotypes. De-Arteaga et al. (De-Arteaga et al., 2019)’s study of semantic representation bias can lead to potential harm.
@@ 2.2. Recommender Systems with Side Information
The adaptation of LLMs to recommendation scenarios has primarily relied on techniques like fine-tuning and in-context learning. Wang et al. (Wang et al., 2023) utilised ChatGPT for session-based recommendations via zero-shot prompting. Agent4Rec (Zhang et al., 2024b) employs LLM-enabled generative agents to simulate users using user side information specifically tailored for recommender systems. Geng et al. introduced a unified framework integrating five recommendation tasks by fine-tuning LLMs. TALLRec (Bao et al., 2023) fine-tunes LLMs for natural language generation tasks via instruction fine-tuning. However, the direct use of LLMs in recommendation tasks faces challenges such as high computational costs and slow inference times. Ren et al. (Ren et al., 2024) used mutual information maximisation to combine LLM knowledge with collaborative relationship modeling. Lu et al. (Lu et al., 2024) employed reinforcement learning to enhance LLMs’ ability to respond to user intent and mitigate formatting errors. Zhang et al. (Zhang et al., 2024a) converted collaborative information from external models into binary sequences for alignment with LLMs.
@@ 3.1. Problem Setting
With an existing recommendation method $\mathcal{R}$, we can obtain the user collaborative representation $\mathbf{\hat{e}}_{u}$ and item collaborative representation $\mathbf{\hat{e}}_{i}$ based on the historical interactions between the user $\mathcal{U}$ and the item $\mathcal{I}$, and then compute the probability that $\mathcal{U}$ and $\mathcal{I}$ will interact with each other in the future. Concurrently, we leverage an LLM to generate user side information $\mathcal{P}_{u}$ and item side information $\mathcal{P}_{i}$, which are then projecting into the feature space via a text embedding model $\mathcal{T}$. An encoder is subsequently utilised to learn the user side representation $\mathbf{\hat{s}}_{u}$ and item side representation $\mathbf{\hat{s}}_{i}$. Finally, these side representations $\mathbf{\hat{s}}_{u}$ and $\mathbf{\hat{s}}_{i}$ are aligned with the collaborative representations $\mathbf{\hat{e}}_{u}$ and $\mathbf{\hat{e}}_{i}$ through contrastive learning, enhancing the overall recommendation performance of the model. However, LLM produces side information often with propensity bias, which has an impact on recommender systems. The aim of this paper is to improve the effectiveness of LLM for recommender systems by mitigating the propensity bias. A summary of the notations used in this paper is shown in Table 1.
@@ 3.2. Graphical Causal Model
Figure 3. Example of causal DAG where $X$, $Y$, and $M$ denote exposure, outcome and mediator variables, respectively. Gray nodes indicate that the variables are set to reference values (e.g., $X$ = $x^{*}$), the half shaded node represents the result affected by the reference value.
@@ 3.3. Causal Effect and Counterfactual Inference
The causal effect of the treatment, e.g. $X$ on the outcome, e.g. $Y$ is defined as the difference between two counterfactual outcomes that would be observed if an individual were exposed to different levels of the exposure of interest (Hicks and Tingley, 2011; Imai et al., 2010; Cheng et al., 2022; Xu et al., 2023). In formal terms, if we denote the outcome as $Y$, and the exposure values of interest as $x$ and $x^{*}$, the causal effect for an individual $j$ is $Y_{j}(x)-Y_{j}(x^{*})$, where $Y_{j}(x)$ is the counterfactual outcome for individual $j$ under exposure level $x$, and $Y_{j}(x^{*})$ is the counterfactual outcome under exposure level $x^{*}$.
The total effect (TE) (Pearl, 2009; Rijnhart et al., 2021) is the overall change in the outcome that would be observed if an individual’s exposure level were changed $x$, taking into account causal paths from the treatment to the outcome, including both direct path and indirect paths via mediators. Mathematically, it is represented as:
(1) $TE=Y_{j}(x,M_{j}(x))-Y_{j}(x^{*},M_{j}(x^{*}))$
@@
The Total Effect is crucial as it provides a complete picture of the impact of an exposure on an outcome, incorporating both the direct effects and the indirect effects mediated through a mediator. The counterfactual results in the mediator model depend not only on the exposure value but also on the mediator value. Counterfactual inference involves hypothetically intervening to change the value of a single variable and imagining the consequences of this change on other variables within the system. As illustrated in Figure 3, the grey nodes indicate that the corresponding variable has been set to a reference value independently of its normal causal mechanisms. Reference states allow the causal graph to be partitioned into sections representing different counterfactual scenarios.
The Pure Natural Direct Effect (PNDE) (Pearl, 2009; Rijnhart et al., 2021) is the effect of the exposure on the outcome that occurs without any influence from the mediator:
@@
When we want to explore the effects through the mediator variable path after removing the direct effect, we can subtract the PNDE from the TE to obtain the total natural indirect effect (TNIE) (Pearl, 2009; Rijnhart et al., 2021):
(3) $TNIE=TE-PNDE=Y_{j}(x,M_{j}(x))-Y_{j}(x,M_{j}(x^{*}))$
where $Y_{j}(x,M_{j}(x))$ is the counterfactual outcome for individual $j$ under exposure level $x$ and the mediator level that would be observed under exposure level $x$, and $Y_{j}(x,M_{j}(x^{*}))$ is the counterfactual outcome under exposure level $x$ but with the mediator level that would be observed under exposure level $x^{*}$.
@@ 4.1. Constructing Side Information
Previous research has demonstrated the importance of side information (Wei et al., 2022; Zhang et al., 2023b; Ren et al., 2024), also known as descriptions of users and items, for collaborative filtering recommender systems to perform effective representation learning. Specifically, user and item edge information that effectively characterises preferences and attractiveness is crucial for modeling latent relationships between users and items.
@@
Compared to simply aggregating sparse profile fields, our approach allows the LLM to leverage its powerful language understanding and commonsense knowledge to generate more informative and coherent side information. The specific formulas are as follows:
@@ 4.2. Spectrum-based Side Information Encoder
The set of parameters for SSE is defined as $\boldsymbol{\theta}=(f,T,\lambda)$. In SSE, we introduce a conditional prior distribution $p_{T,\lambda}(Z\mid m)$, where $Z$ denotes the latent variables and $m$ represents additional observed variables. In our SSE framework, we use the spectrum of the historical interaction information. We have added a small noise to the spectrum to avoid over-fitting during the learning process as follows:

### v2

@@ Abstract.
*Corresponding authors
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: noneccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems are essential tools used on many platforms to help users navigate the vast amount of available information and find items that match their preferences (Gao et al., 2023; Wang et al., 2018; Wu et al., 2023). Due to the information overload problem (Zhang et al., 2017; He et al., 2020; Deng et al., 2022), collaborative filtering-based recommender systems have become widely used. The main research direction in this field focuses on improving the performance of these models by increasing the effectiveness of user and item representations (Zhang et al., 2019). To achieve this, many models have been proposed to mine useful information to enhance these representations. Graph neural networks play an important role in this endeavor (Wei et al., 2022b; Sun et al., 2022; Wu et al., 2024b; Zhang et al., 2024d). Recently, graph neural network-based recommender systems modeling user preferences based on historical interactions have achieved promising results (Wu et al., 2022a; Liu et al., 2022; Wei et al., 2022a; Bao et al., 2025; Guan et al., 2024; Liu et al., 2024a). Nevertheless, collaborative filtering models still face the problem of data sparsity (Cai et al., 2023; Yu et al., 2023a).
Researchers have argued that relying solely on interaction information may ignore other valuable data (Gao et al., 2023; Wang et al., 2020), such as rich information related to users and items in the real world. The lack of additional information can lead to a less informative representation learning. Previous research has demonstrated the importance of side information, also known as user and item descriptions, for collaborative filtering recommender systems to perform effective representation learning (Wei et al., 2022b; Zhang et al., 2023b; Ren et al., 2024). However, these approaches tend to be based on domain-specific knowledge and are not sufficiently generalised for recommendation tasks that they do not know much about (Peng et al., 2023).
@@
Recently, Large Language Models (LLMs) have garnered attention due to their rich knowledge and powerful reasoning capabilities, with numerous studies (Lu et al., 2024b; Wei et al., 2024; Rong et al., 2024) showcasing their superior performance. These LLMs, equipped with the ability to process vast amounts of text data, offer promising opportunities for enhancing traditional recommender systems. However, directly deploying off-the-shelf LLMs for recommendation tasks presents several challenges (Wei et al., 2024; Ren et al., 2024; Zhang et al., 2024a; Lu et al., 2024a). A key issue lies in the inherent mismatch between the generalisation capabilities of LLMs and the specific demands of personalised recommendation, particularly when it comes to capturing the nuanced preferences and behaviors of individual users. To bridge this gap, researchers have attempted to leverage LLMs to model human-like descriptions of users and items, thereby obtaining richer side information. The enriched data can then be aligned with collaborative representations learned from users’ historical interactions via techniques such as contrastive learning (Ren et al., 2024), representing a promising direction for integrating the reasoning capabilities of LLMs with the precision required for personalised recommendations.
However, existing research has pointed out that the integration of LLM may bring new issues and challenges (Ma, 2024; Wu et al., 2024a; Liu et al., 2024b). Specifically, the inherent biases in LLM training datasets (Jiang et al., 2024) inevitably introduce propensity bias, which may generate inaccurate or irrelevant content. Specifically, LLM may exhibit biases toward certain cultures and viewpoints (Tao et al., 2023; Yu et al., 2023b; Sivaprasad et al., 2024), but not all propensity biases are harmful. Real-world knowledge exists to favour certain items possibly because they possess good intrinsic properties, and blindly removing this part of the propensity bias (named user/item propensity bias in this paper) would make the model lose such real-world knowledge. However, the propensity bias which is not related to users/items (named user/item unrelated propensity bias in this paper) can result in overly generalized representation, ignoring user and item specific details. As shown in Figure 1, a user with ID 5596 in the Amazon dataset (He and McAuley, 2016) has previously read a range of thrilling and exciting books, yet the LLM overlooks this in its reasoning and generalises the user’s preference as simply “entertaining”.
@@
Furthermore, the unrelated propensity bias not only impacts recommendation accuracy but also contributes to the homogenisation of LLM-generated representations (Barbero et al., 2024), which increases the risk of dimensional collapse. When using LLMs (Guo et al., 2024) or contrastive learning (Jing et al., 2022), side representations are more likely to be restricted to a certain portion of the low-dimensional subspace, futher limiting their representational capability (Zhang et al., 2023c). As shown in Figure. 2, when using LightGCN as the backbone method to align LLMs for recommendation through contrastive learning, the model exhibits dimensional collapse when side information is encoded using MLP. This occurs because the side information repsentation generated by the LLM lacks structural information, making it difficult to align with collaborative information derived from historical interactions, thereby exacerbating the risk of dimensional collapse.
@@ 2.1. Bias in Artificial Intelligence
With the widespread use of artificial intelligence in real life (Deng et al., 2016, 2021, 2023; Zhang et al., 2022; Zhang and Li, 2023; Zhang, 2022), bias in artificial intelligence has attracted a great deal of attention (Zhang et al., 2023a; Ling et al., 2024). In a causal inference perspective (Wang et al., 2023a; Li and Yao, 2024), bias is often due to spurious correlations caused by confounders (Zhang et al., 2025). For example, Buolamwini and Gebru (Buolamwini and Gebru, 2018) showed that business gender classification systems used in various fields such as marketing, entertainment and healthcare have different representations of users’ skin colour and gender. Zhang et al. (Zhang et al., 2024c) demonstrated that the interaction structure in social networks can also contain preference information. It follows that artificial intelligence systems may display biases against specific groups embedded in the data or even amplify human biases. Natural Language Processing (NLP) systems are susceptible to bias as a result of the training corpus. Bolukbasi et al. (Bolukbasi et al., 2016) and Caliskan et al. (Caliskan et al., 2017) first found the presence of human-like preferences in word embedding models. Zhao et al. (Zhao et al., 2018) found a correlation between gender pronouns and stereotypes. De-Arteaga et al. (De-Arteaga et al., 2019)’s study of semantic representation bias can lead to potential harm.
@@ 2.2. Recommender Systems with Side Information
The adaptation of LLMs to recommendation scenarios has primarily relied on techniques like fine-tuning and in-context learning. Wang et al. (Wang et al., 2023b) utilised ChatGPT for session-based recommendations via zero-shot prompting. Agent4Rec (Zhang et al., 2024b) employs LLM-enabled generative agents to simulate users using user side information specifically tailored for recommender systems. Geng et al. introduced a unified framework integrating five recommendation tasks by fine-tuning LLMs. TALLRec (Bao et al., 2023) fine-tunes LLMs for natural language generation tasks via instruction fine-tuning. However, the direct use of LLMs in recommendation tasks faces challenges such as high computational costs and slow inference times. Ren et al. (Ren et al., 2024) used mutual information maximisation to combine LLM knowledge with collaborative relationship modeling. Lu et al. (Lu et al., 2024a) employed reinforcement learning to enhance LLMs’ ability to respond to user intent and mitigate formatting errors. Zhang et al. (Zhang et al., 2024a) converted collaborative information from external models into binary sequences for alignment with LLMs.
@@ 3.1. Problem Setting
With an existing recommendation method $\mathcal{R}$, we can obtain the user collaborative representation $\mathbf{\hat{e}}_{u}$ and item collaborative representation $\mathbf{\hat{e}}_{i}$ based on the historical interactions between the user $\mathcal{U}$ and the item $\mathcal{I}$, and then compute the probability that $\mathcal{U}$ and $\mathcal{I}$ will interact with each other in the future. Concurrently, we leverage an LLM to generate user side information $\mathcal{P}_{u}$ and item side information $\mathcal{P}_{i}$, which are then projecting into the feature space via a text embedding model $\mathcal{T}$. An encoder is subsequently utilised to learn the user side representation $\mathbf{\hat{s}}_{u}$ and item side representation $\mathbf{\hat{s}}_{i}$. Finally, these side representations $\mathbf{\hat{s}}_{u}$ and $\mathbf{\hat{s}}_{i}$ are aligned with the collaborative representations $\mathbf{\hat{e}}_{u}$ and $\mathbf{\hat{e}}_{i}$ through contrastive learning, enhancing the overall recommendation performance of the model. However, LLM produces side information often with propensity bias, which has an impact on recommender systems. The aim of this paper is to improve the effectiveness of LLM for recommender systems by mitigating the unrelated propensity bias. A summary of the notations used in this paper is shown in Table 1.
@@ 3.2. Graphical Causal Model
Figure 3. Example of causal DAG where $X$, $Y$, and $M$ denote exposure, outcome and mediator variables, respectively. Gray nodes indicate that the variables are set to reference values (e.g., $X$ = $x^{*}$), the half shaded node represents the result affected by the reference value. In this figure, (a) represents the causal graph for the factual scenario, (b) shows the causal graph where X is set to the reference state, and (c) illustrates the causal graph for the counterfactual scenario, where X is set to different values.
@@ 3.3. Causal Effect and Counterfactual Inference
The causal effect of the exposure, e.g. $X$ on the outcome, e.g. $Y$ is defined as the difference between two counterfactual outcomes that would be observed if an individual were exposed to different levels of the exposure of interest (Hicks and Tingley, 2011; Imai et al., 2010; Cheng et al., 2022; Xu et al., 2023). In formal terms, if we denote the outcome as $Y$, and the exposure values of interest as $x$ and $x^{*}$, the causal effect for an individual $j$ is $Y_{j}(x)-Y_{j}(x^{*})$, where $Y_{j}(x)$ is the counterfactual outcome for individual $j$ under exposure level $x$, and $Y_{j}(x^{*})$ is the counterfactual outcome under exposure level $x^{*}$.
To ensure a valid causal interpretation of the Pure Natural Direct Effect (PNDE) and Total Natural Indirect Effect (TNIE) (Pearl, 2009; Rijnhart et al., 2021) at the population-average level, the following four assumptions must hold (Pearl, 2009; Rijnhart et al., 2021; Cheng et al., 2022; Imbens and Rubin, 2015; Xu et al., 2023):
  • No unmeasured confounding of the exposure-outcome effect: All confounders that affect both the exposure $X$ and the outcome $Y$ must be observed and appropriately adjusted for. This ensures that the direct effect of $X$ on $Y$ is not confounded by unmeasured variables.
  • No unmeasured confounding of the mediator-outcome effect: All variables influencing the mediator $M$ and the outcome $Y$ must be measured. This ensures that the indirect effect of $X$ on $Y$ via $M$ is not biased by hidden confounders.
  • No unmeasured confounding of the exposure-mediator effect: Every variable that confounds the relationship between the exposure $X$ and the mediator $M$ must be observed and accounted for. This ensures that the causal effect of $X$ on $M$ is not confounded by unmeasured variables.
  • No confounders of the mediator-outcome effect that are affected by the exposure: This assumption requires that no variables confounding the mediator-outcome relationship $M\rightarrow Y$ are themselves influenced by the exposure $X$.
The total effect (TE) (Pearl, 2009; Rijnhart et al., 2021) is the overall change in the outcome that would be observed if an individual’s exposure level were changed $x$, taking into account causal paths from the exposure to the outcome, including both direct path and indirect paths via mediators. Mathematically, it is represented as:
(1) $TE=Y_{j}(x,M_{j}(x))-Y_{j}(x^{*},M_{j}(x^{*})),$
@@
The TE is crucial as it provides a complete picture of the impact of an exposure on an outcome, incorporating both the direct effects and the indirect effects mediated through a mediator. The counterfactual results in the mediator model depend not only on the exposure value but also on the mediator value. Counterfactual inference involves hypothetically intervening to change the value of a single variable and imagining the consequences of this change on other variables within the system. As illustrated in Figure 3 (b), the grey nodes indicate that the corresponding variable has been set to a reference value independently of the facts. As shown in Figure 3 (c), the introduction of reference states helps to divide the causal graph into different parts, i.e. to realize a counterfactual scenario. This scenario is termed a ‘counterfactual scenario’ because it entails the simultaneous consideration of both a fact state, denoted by $X=x$, and a reference state, represented by $X=x^{*}$.
The PNDE is the effect of the exposure on the outcome that occurs without any influence from the mediator:
@@
When we want to explore the effects through the mediator variable path after removing the direct effect, we can subtract the PNDE from the TE to obtain the TNIE:
(3) $TNIE=TE-PNDE=Y_{j}(x,M_{j}(x))-Y_{j}(x,M_{j}(x^{*})),$
where $Y_{j}(x,M_{j}(x))$ is the potential outcome for individual $j$ under exposure level $x$ and the mediator level that would be observed under exposure level $x$, and $Y_{j}(x,M_{j}(x^{*}))$ is the counterfactual potential outcome under exposure level $x$ but with the mediator level that would be observed under exposure level $x^{*}$.
@@ 4.1. Constructing Side Information
Previous research has demonstrated the importance of side information (Wei et al., 2022b; Zhang et al., 2023b; Ren et al., 2024), also known as descriptions of users and items, for collaborative filtering recommender systems to perform effective representation learning. Specifically, user and item edge information that effectively characterises preferences and attractiveness is crucial for modeling latent relationships between users and items.
@@
Compared to simply aggregating sparse profile fields, our approach allows the LLM to leverage its powerful language understanding and commonsense knowledge to generate more informative and coherent side information. After obtaining the textual description of the side information given by the LLM, it is necessary to transform the text into embeddings using the following formulas:
@@ 4.2. Spectrum-based Side Information Encoder
The set of parameters for SSE is defined as $\boldsymbol{\theta}=(f,T,\lambda)$. In SSE, we introduce a conditional prior distribution $p_{T,\lambda}(Z\mid M)$, where $Z$ denotes the latent variables and $M$ represents additional observed variables. In our SSE framework, the spectrum of historical interaction information is derived through singular value decomposition (SVD). However, users may occasionally click by mistake when performing an action, and overfitting historical interaction data may result in overfitting, which may not accurately capture user preferences. To avoid overfitting and increase the generalisation capabilities of the model, we add a small amount of noise to the spectrum: